# Vision-driven Automated Mobile GUI Testing via Multimodal Large Language Model

### v1

@@ IV-B Baselines
For intra-page non-crash bug detection, the techniques are mainly for GUI display-related issue detection. OwlEyes [12] and NightHawk [13] use deep learning to detect UI display issues (e.g., overlapping texts, component occlusion). DVermin [15] detects scaling issues by comparing GUI pages under different page settings. DiffDroid [14] compares the GUI pages of an app on two different devices to find compatibility issues. The input of these techniques are screenshots/visual hierarchy files and need to be integrated with automated testing approaches to fetch the required information during the exploration. We utilize four popular automated GUI testing tools, i.e., model-based tools APE [9] and Fastbot [34], learning-based tool Humanoid [35] and LLM-based GPTDroid [24].
@@
For inter-page non-crash bug detection, we choose 6 baselines from different perspectives. iFixdataloss [36] is the most recent work of finding data loss issues. SetDroid [37] uses metamorphic testing to find system setting-related functional bugs. Genie [38] uses the metamorphic relation of independent view property to find non-crash functional bugs, while Odin [29] uses the implicit knowledge of “bugs as deviant behavior” to find non-crash functional bugs. Android Lint [39] is a popular static analysis tool for finding different issues in Android apps based on predefined rules. LiveDroid [40] is also a static analysis tool for detecting data loss issues.
IV-C Experimental Setup
@@
We deploy the baselines and our approach on a 64-bit Ubuntu 18.04 machine (64 cores, AMD CPU, 2GB RAM) and evaluate them on Google Android 7.1 emulators. Following common practice [7, 8, 10], we registered separate accounts for each bug that requires login and wrote the login scripts, and during testing reset the account data before each run to avoid possible interference. To ensure fair and reasonable use of resources, we set up the running time of each tool in one app to 50 minutes, which is widely used in other GUI testing studies [8, 10, 41, 34] We run each tool 3 times and take the average value as the result to mitigate potential bias [8, 10, 41, 34].
@@
VisionDroid 0.57 0.55 109 55 0.50 0.42 0.65 0.56 177 127 0.72 0.51 0.61 0.59 38 29 0.76 0.64
w/o Alignment 0.44 0.39 97 32 0.33 0.24 0.49 0.48 133 75 0.56 0.30 0.47 0.43 29 18 0.62 0.40
w/o History 0.41 0.35 91 28 0.31 0.21 0.47 0.44 128 63 0.49 0.26 0.44 0.31 27 15 0.56 0.33
w/o Community 0.57 0.55 90 17 0.19 0.13 0.65 0.56 149 39 0.26 0.16 0.61 0.59 42 7 0.17 0.16
w/o Example 0.57 0.55 89 18 0.20 0.14 0.65 0.56 151 41 0.27 0.17 0.61 0.59 39 8 0.21 0.18
IV-D Evaluation Metrics
@@
We measure the performance from coverage and bug detection, respectively corresponding to the explorer and detector in VisionDroid. For coverage, we obtain the activity coverage and code coverage [42, 43, 44, 10, 45, 35], in which treat the activities defined in the AndroidManifest.xml file of an Android app as the whole set of activities [46, 47, 7].
@@
For bug detection, following existing studies [38, 29], we report the detected bugs by VisionDroid, true bugs among the detected bugs. Since we know the entire set of bugs, we also calculate the precision and recall of the bug detection.
V Results and Analysis
V-A Coverage and Bug Detection Performance (RQ1)
@@
We present the coverage and bug detection performance of VisionDroid and baselines, in terms of the baseline dataset, GitHub dataset and the bug injection dataset in Table I.
@@
We first present the coverage of VisionDroid, which can cover far more activities and code than the baselines. In three datasets, the average activity coverage is between 57% and 61%, and the average code coverage is between 55% and 59%. It achieves between 21% (0.57 vs. 0.47) to 33% (0.65 vs. 0.49) activity coverage and between 27% (0.56 vs. 0.44) and 43% (0.59 vs. 0.41) code coverage higher even compared with the best baseline (GPTDroid) across three datasets. This indicates the effectiveness of VisionDroid in covering more activities and codes, thus bringing higher confidence to app quality, potentially uncovering more bugs.
@@
The average precision and recall of VisionDroid in three datasets achieve 50%-76%, and 42%-64%, respectively. It achieves between 8% (0.72 vs. 0.67) and 230% (0.76 vs. 0.23) precision and between 27% (0.56 vs. 0.44) and 43% (0.59 vs. 0.41) higher even compared with the best baseline across three datasets. Please note that although some baseline methods can achieve higher precision, they detect fewer bugs. In terms of the number of detected bugs, VisionDroid can still achieve a higher detection rate. We also compare the similarities and differences of the bugs between the best baselines (Odin and NightHawk) and our approach, and the results show that all bugs detected by them are also detected by VisionDroid. This indicates the effectiveness of VisionDroid in detecting bugs and helps to ensure app quality.
@@
The advantages of VisionDroid are due to three aspects. First, VisionDroid designs the vision prompt which can better capture the GUI context, while the baselines mainly rely on the view hierarchy file in which the widgets might have meaningless “text” or “resource-id”, hindering them from effectively understanding the GUI pages. Second, VisionDroid plans the exploration path both considering the semantics of a single GUI page and the exploration history which can enable the function-aware exploration, while the baselines can hardly understand the inherent business logic of the app, resulting in aimless exploration. Third, VisionDroid detects the bugs by understanding the semantics of a sequence of GUI pages, while the baselines can only detect specific types of bugs or have low detection accuracy.
Fig. 7 demonstrates an example of VisionDroid’s detected bug. This bug is that after deleting a single expense on the budget information page, all original expense records disappear when viewing the expense page. We can see that VisionDroid is adept at determining whether the functional logic of the app is abnormal based on the semantics of test path and GUI information.
V-B Ablation Study (RQ2)
V-B1 Contribution of Sub-modules
Table II further demonstrates the performance of VisionDroid and its 4 variants. We remove each sub-module of the VisionDroid in Fig. 3 separately, i.e., Alignment text with image, Function-based testing history construction, Function-driven community detection and Example bugs retrieval. For removing the Alignment text with image, we directly input the unmarked screenshot. For removing function-based testing history construction, we delete the history information. For removing the community detection, we randomly segment the test sequence and for removing retrieved examples, we use the random selection.
The experiment results (Table II) demonstrate that removing any of the sub-modules would result in a noticeable performance decline, indicating the necessity and effectiveness of the designed sub-modules. Removing the community detection (w/o-Community) has the greatest impact on the performance, reducing the bug detection performance by 347% (0.17 vs. 0.76) precision and 300% (0.16 vs. 0.64).
We also notice that, when removing the alignment text with image (w/o-Alignment) and function-based testing history construction(w/o-History), the activity coverage has an impact, reducing the activity coverage by 39% (0.41 vs. 0.57) and 57% (0.35 vs. 0.55). This further indicates that the alignment text with image can help MLLM understand the structure and semantic information of GUI pages and make reasonable judgments. The history can help retain knowledge during testing and gain global viewpoints to reach uncovered areas.
@@ V-B1 Contribution of Sub-modules
Besides, randomly selecting examples (w/o-retriExample) would also largely influence the performance, and decrease the bug detection performance by 262% (0.21 vs. 0.76) precision and 256% (0.18 vs. 0.64) recall. This indicates that by providing similar examples, the MLLM can quickly think out what should the no-crash bugs look like, which further indicates the demonstration can facilitate the MLLM in producing the required output.
V-B2 Influence of Different Number of Examples
@@
Figure 9: Different number of examples (RQ2)
V-C Usefulness Evaluation (RQ3)
For the 105 apps, VisionDroid detects 83 non-crash bugs in 53 apps, of which 37 bugs were never discovered before. Furthermore, only 9 of these new bugs are detected by Genie [38] and Odin [29]. We submitted these 37 bugs to developers, and 19 of them have been fixed/confirmed so far (10 fixed and 9 confirmed), while the remaining are still pending (none of them is rejected). This further indicates the effectiveness of our proposed VisionDroid in bug detection. Due to space limit, Table III presents these fixed/confirmed bugs, and the full lists can be found on our website1. We further analyze the details of these bugs, and 21 of them involve intra-page bugs and 16 of them involve inter-page bugs.
TABLE III: Confirmed or fixed bugs. (RQ3)
Id APP Name Category Download Status
1 MOMPA Music 500M+ fixed
2 Calendar Tool 100M+ fixed
3 DigiCal Tool 100M+ fixed
4 Expensify Finance 100M+ fixed
5 TrackWallet Finance 10M+ confirmed
6 Roadie Travel 10M+ confirmed
7 ISAY Commun 10M+ fixed
8 EasyBudget Finance 10M+ confirmed
9 Iplan Travel 10M+ fixed
10 MediaFire Product 5M+ confirmed
11 SmartGuide Navig 500K+ fixed
12 MMDR Utilities 500K+ confirmed
13 Sygic Travel 500K+ fixed
14 Fair Health 500K+ confirmed
15 ClassySha Tool 500K+ fixed
16 Linphone Commun 50K+ confirmed
17 Rocket Finance 50K+ confirmed
18 Monefy Finance 50K+ fixed
19 Origin Finance 50K+ confirmed
@@ VI-A Automated GUI Testing
To ensure the quality of mobile apps, many researchers study the automatic generation of large-scale test scripts to test apps [48]. Since Android apps are event-based [49, 50, 51, 52], the most common automated testing methods are model-based automated GUI testing methods [2, 3, 4, 5, 6, 7, 8, 9, 10], design corresponding models through the analysis of the apps. Due to the lack of consideration for the semantic information of the app’s GUI pages, the coverage of model-based methods is still low. Researchers further proposed human-like testing strategies and designed learning-based automated GUI testing methods [35, 46]. GPTDroid [24] used GPT-3.5 to generate the testing script, aiming at proposing a more effective approach to generate human-like actions for thoroughly and more effectively testing the app. However, these methods require first converting the GUI visual information into text, which can cause significant visual information loss during the process. Moreover, a large number of app widgets and pages are generated through dynamic rendering, and the attribute information of image widgets will be lost when obtaining view hierarchy files. As a result, the coverage of automated testing is low, and a large number of widgets and operations are missed. This study employs the MLLM for the GUI testing, which is more effective through aligning the GUI screenshot with text information for the vision prompt.
@@ VII Conclusion
With the advancement of LLM, there have been tremendous applications of LLMs in software engineering-related tasks such as program repair and test case generation with great performance improvement. Subsequently, MLLMs emerge and are applied in image-text reasoning, video generation, etc., further expanding our understanding of the capabilities of LLMs. This paper proposes VisionDroid, the pioneering work using MLLMs for vision-based mobile app GUI testing, and the experiment demonstrates its effectiveness and usefulness. This work contributes as a stepping stone for future exploration in software engineering, offering insights and guidance on the potential applications of MLLM. In the future, we will further improve this approach for better performance. Meanwhile, beyond Android apps, we will customize and apply our approach to other platforms such as iOS, web, desktop and even embedded systems for software testing.

### v2

@@ III-B Baselines
There are 4 mobile app display issue detection techniques. OwlEyes [11] and NightHawk [12] use deep learning to detect UI display issues (e.g., overlapping texts, component occlusion). DVermin [31] detects scaling issues by comparing GUI pages under different page settings. DiffDroid [32] compares the GUI pages of an app on two different devices to find compatibility issues. The input of these techniques are screenshots/visual hierarchy files and need to be integrated with automated testing approaches to fetch the required information during the exploration. We respectively pair them with five popular automated GUI testing tools, i.e., random-based tool Monkey [33], model-based tools APE [8] and Fastbot [34], learning-based tool Humanoid [15] and LLM-based GPTDroid [21].
@@
We utilize 6 baselines of logical bug detection for mobile apps. iFixdataloss [35] is the most recent work of finding data loss issues. SetDroid [36] uses metamorphic testing to find system setting-related functional bugs. Genie [37] uses the metamorphic relation of independent view property to find non-crash functional bugs, while Odin [26] uses the implicit knowledge of “bugs as deviant behavior” to find non-crash functional bugs. Android Lint [38] is a popular static analysis tool for finding different issues in Android apps based on predefined rules. LiveDroid [39] is also a static analysis tool for detecting data loss issues.
We also employ 2 MLLM-based methods as our baselines. AppAgent [40] is a novel MLLM-based framework that enables agents to interact with apps through a simplified action space, mimicking human-like interactions such as tapping and swiping. We modify AppAgent by adding a query in its prompt after each test interaction to check for any detected bugs on the page. Additionally, we include GPT-4 as another baseline. We provided GPT-4 with annotated screenshots of the app. GPT-4 was then tasked with generating actions based on these screenshots and detecting any potential bugs.
III-C Experimental Setup
@@
We deploy the baselines and our approach on a 64-bit machine and evaluate them on Google Android emulators. Following common practice [6, 7, 9], we registered separate accounts for each bug that requires login and wrote the login scripts, and during testing reset the account data before each run to avoid possible interference. To ensure fair and reasonable use of resources, we set up the running time of each tool in one app to 50 minutes, which is widely used in other GUI testing studies [7, 9, 16, 34]. We run each tool 3 times and take the average value as the result to mitigate potential bias [7, 9, 16, 34].
@@
Trident 0.57 0.55 109 55 0.50 0.42 0.65 0.62 278 201 0.72 0.52 0.61 0.59 67 47 0.70 0.65
w/o Screenshot annotation 0.41 0.48 55 27 0.49 0.21 0.41 0.39 101 49 0.49 0.13 0.38 0.36 39 15 0.38 0.21
w/o GUI information 0.49 0.47 69 33 0.48 0.25 0.52 0.50 222 129 0.58 0.33 0.49 0.45 52 31 0.60 0.43
w/o Testing history 0.39 0.37 43 21 0.49 0.16 0.42 0.39 131 53 0.40 0.14 0.44 0.42 43 19 0.44 0.26
w/o Optimization 0.47 0.45 39 16 0.41 0.12 0.44 0.42 87 38 0.44 0.10 0.42 0.41 42 17 0.40 0.24
w/o Image Legend 0.38 0.36 27 12 0.44 0.09 0.39 0.38 78 31 0.40 0.08 0.42 0.41 39 15 0.38 0.21
w/o Test Recoding 0.42 0.39 38 18 0.47 0.14 0.48 0.46 98 51 0.52 0.13 0.43 0.42 41 19 0.46 0.26
w/o Bug example 0.53 0.52 47 9 0.19 0.07 0.59 0.58 51 17 0.33 0.04 0.54 0.52 38 5 0.13 0.07
w/o Function COT 0.51 0.49 54 11 0.20 0.08 0.58 0.55 63 23 0.37 0.06 0.55 0.53 39 7 0.18 0.10
III-D Evaluation Metrics
@@
We measure the performance from coverage and bug detection, respectively corresponding to the explorer and detector in Trident. For coverage, we obtain the activity coverage and code coverage [41, 42, 43, 9, 44, 15], in which we treat the activities defined in the AndroidManifest.xml file of an Android app as the whole set of activities [17, 45, 6].
@@
For bug detection, following existing studies [37, 26], we report the detected bugs by Trident, true bugs among the detected bugs. Since we know the entire set of bugs, we also calculate the precision and recall of the bug detection.
IV Results and Analysis
IV-A Coverage and Bug Detection Performance (RQ1)
@@
We present the coverage and bug detection performance of Trident and baselines, in terms of the baseline dataset, GitHub dataset and the bug injection dataset in Table I.
@@
We first present the coverage of Trident, which can cover far more activities and code than the baselines. In three datasets, the average activity coverage is between 57% and 65%, and the average code coverage is between 55% and 62%. It achieves between 21% (0.57 vs. 0.47) to 27% (0.65 vs. 0.51) activity coverage, and between 28% (0.55 vs. 0.43) and 32% (0.62 vs. 0.47) code coverage higher even compared with the best baseline (GPTDroid) across three datasets. This indicates the effectiveness of Trident in covering more activities and codes, thus bringing higher confidence to app quality, potentially uncovering more bugs.
@@
The average precision and recall of Trident in three datasets achieve 50%-72%, and 42%-65%, respectively. It achieves between 14% (0.72 vs. 0.63) and 112% (0.70 vs. 0.33) precision and between 108% (0.52 vs. 0.25) and 147% (0.42 vs. 0.17) recall higher even compared with the best baseline across three datasets. Please note that although some baseline methods can achieve higher precision, they detect fewer bugs. In terms of the number of detected bugs, Trident can still achieve a higher detection rate. We also compare the similarities and differences of the bugs between the best baselines (Odin and NightHawk) and our approach, and the results show that all bugs detected by them are also detected by Trident. This indicates the effectiveness of Trident in detecting bugs and helps to ensure app quality.
@@
The advantages of Trident are due to three aspects. First, Trident designs the vision prompt which can better capture the GUI context, while the baselines mainly rely on the view hierarchy file in which the widgets might have meaningless “text” or “resource-id”, hindering them from effectively understanding the GUI pages. Second, Trident plans the exploration path both considering the semantics of a single GUI page and the exploration history which can enable the function-aware exploration, while the baselines can hardly understand the inherent business logic of the app, resulting in aimless exploration. Third, Trident detects the bugs by understanding the semantics of a sequence of GUI pages, while the baselines can only detect specific types of bugs or have low detection accuracy.
Fig. 1 and Fig. 8 show some examples of bugs detected by Trident. These bugs belong to different types and are not detected by the baseline method. In Fig. 8 (a), when the user adds income, enter \$520. After returning to the main page, the income was repeatedly displayed in the list, resulting in the same income being incorrectly recorded twice. In Fig. 8 (b), the user clicks on the “Devices” tab and then clicks on the “Notifications”. As a result, the app returns to the main page instead of the correct notification page. We can see that Trident is adept at determining whether the functional logic of the app is abnormal based on the semantics of the test path and GUI information.
Figure 8: Example of bugs detected by Trident
IV-B Ablation Study (RQ2)
IV-B1 Contribution of Sub-modules
Table II further demonstrates the performance of Trident and its 8 variants. We remove each sub-module of the Trident in Fig. 2 separately, i.e., Screenshot annotation, GUI information, Testing history, Optimization, Image Legend, Test Recoding, Bug example and Function COT. For removing the Screenshot annotation, We adopt the direct annotation method shown in Fig. 3 (a). For removing GUI information, Testing history, and Optimization, we separately delete three corresponding parts from the prompt in Fig. 4. For removing the Image Legend, we delete the image legend in three prompts. For removing the Test Recoding, we delete it from the prompt in Fig. 6. For removing the Bug example and Function COT, we delete them from the prompt in Fig. 7.
@@ IV-B1 Contribution of Sub-modules
The experiment results (Table II) demonstrate that removing any of the sub-modules would result in a noticeable performance decline, indicating the necessity and effectiveness of the designed sub-modules. Removing the bugs example (w/o-Bug example) has the greatest impact on the performance, reducing the bug detection performance by 54% (0.33 vs. 0.72) precision and 92% (0.04 vs. 0.52) recall. This indicates that by providing similar examples, the MLLM can quickly think out what should the non-crash bugs look like, which further indicates the demonstration can facilitate the MLLM in producing the required output.
We also notice that, the w/o-Screenshot annotation, w/o-Testing history and w/o-Image Legend, the activity coverage has an impact, reducing the activity coverage by 37% (0.41 vs. 0.65), 35% (0.42 vs. 0.65) and 40% (0.39 vs. 0.65). This further indicates that the alignment of text with images can help MLLM understand the structure and semantic information of GUI pages and make reasonable judgments. The history can help retain knowledge during testing and gain global viewpoints to reach uncovered areas.
Besides, (w/o-Function COT) would also largely influence the performance, and decrease the bug detection performance by 56% (0.37 vs. 0.72) precision and 88% (0.06 vs. 0.52) recall. This indicates that through COT, MLLM can infer the possible path of the functionality based on its description, and then judge the testing path, improving MLLM’s understanding of the functionality and bug detection performance.
Figure 9: Different number of examples (RQ2)
IV-B2 Influence of Different Number of Examples
@@
TABLE III: Information of the fixed bugs. (RQ3)
Id APP Name Download Category Version
1 Waze 500M+ map 4.103
2 MEGA 100M+ productivity 13.5
3 Swiggy 100M+ food 4.61
4 AlfredCamera 50M+ house 2024.2
5 LG ThinQ 50M+ lifestyle 5.0.2
6 SmartNews 50M+ news 24.7.55
7 Da Fit 50M+ sport 2.7.7
8 Blinkit 50M+ food 16.11.0
9 Bose 5M+ music 10.2.4
10 TomTom 5M+ map 9.51
11 ExpressPlus 5M+ medical 4.12.0
12 Property 5M+ house 5.267.0
13 Plant Parent 5M+ lifestyle 1.76
14 NOS 5M+ news 202405
15 CommBank 10M+ finance 5.11.0
16 AIMP 10M+ music 4.1
17 Wise 10M+ finance 8.71
18 Smart Life 10M+ lifestyle $5.1$
19 Zepp 10M+ sport 8.11
20 Hungerstation 10M+ food 8.0.1
21 Lumosity 10M+ education 2024.02
22 Canvas Student 10M+ education 7.5.0
23 Khan Academy 10M+ education 8.1.1
24 Healthengine 1M+ health 10.1.11
25 Virgin Australia 1M+ travel 2.33.0
26 Opal Travel 1M+ travel 9.8.2
27 NAB Banking 1M+ finance 9.153.0
28 Flybuys 1M+ shopping 24.7.1
29 EverydayRewards 1M+ shopping 24.11.0
30 Woolworths 1M+ shopping 24.14.0
31 Home Assistant 1M+ house 2024.5
IV-C Usefulness Evaluation (RQ3)
For the 187 apps, Trident detects 102 non-crash bugs in 74 apps, of which 43 bugs in 42 apps are never discovered before. Furthermore, only 9 of these new bugs are detected by Genie [37] and Odin [26]. We submitted these 43 bugs to developers, and all of them have been fixed/confirmed so far (31 fixed and 12 confirmed, none of them rejected). This further indicates the effectiveness of our proposed Trident in bug detection. Due to space limit, Table III presents a part of fixed/confirmed bugs, and the full lists can be found on our website1. We also collect the feedback from developers regarding the bugs we submitted.“This bug is crucial because it affects a core functionality that users frequently interact with.”, “We’re aware that this bug is hindering users’ ability. It’s a high priority for us to address this issue.”. These feedbacks also demonstrate the importance of Trident in detecting non-crash bugs.
V Discussion
This Section further discusses the generality of Trident across platforms. Almost the existing studies of non-crash bug detection [35, 36, 37, 26, 39] are designed for a specific platform (Android), which limits its applicability in real-world practice. In comparison, the primary idea of Trident is to explore the app and detect the non-crash functional bugs with a visual understanding of MLLM. Since the screenshots from different platforms (e.g., Android, iOS, Harmony OS) exert almost no difference, Trident can be generalized for App exploration and bug detection in other platforms. For platforms like iOS and Harmony OS, where text extraction might not be as reliable or feasible, we adapt Trident by deleting the textual GUI information from the prompts, and using the remaining modules of Trident to guide the app exploration and bug detection process.
TABLE IV: Cross platform performance of Trident
Platform Code coverage Report bug Real bug Precision
IOS 0.47 15 9 0.60
Harmony 0.57 19 13 0.69
We conduct a small-scale experiment for the popular platform iOS and Harmony OS, and collect 10 open-source apps for testing on each platform. As shown in Table IV, results show that the average code coverage is 0.47 in IOS and 0.57 in Harmony OS. The bug detection precision in IOS is 0.60 and the precision in Harmony OS is 0.60. This is comparable to the performance in Android, and further demonstrates the generality of Trident. We will conduct more thorough experiments in the future.
@@ VI-A Automated GUI Testing
To ensure the quality of mobile apps, many researchers study the automatic generation of large-scale test scripts to test apps [46]. Since Android apps are event-based [47, 48, 49, 50], the most common automated testing methods are model-based automated GUI testing methods [1, 2, 3, 4, 5, 6, 7, 8, 9], design corresponding models through the analysis of the apps. Due to the lack of consideration for the semantic information of the app’s GUI pages, the coverage of model-based methods is still low. Researchers further proposed human-like testing strategies and designed learning-based automated GUI testing methods [15, 17]. GPTDroid [21] used GPT-3.5 to generate the testing script, aiming at proposing a more effective approach to generate human-like actions for testing the app. AppAgent [40] is a novel LLM-based multimodal agent framework, that enables the agent to operate apps through a simplified action space. Qtypist [51] and AudoDroid [52] combine LLM and traditional GUI testing tools during the exploration process. However, these methods require first converting the GUI visual information into text, which can cause significant visual information loss during the process. Moreover, a large number of app widgets and pages are generated through dynamic rendering, and the attribute information of image widgets will be lost when obtaining view hierarchy files. As a result, the coverage of automated testing is low, and many widgets and operations are missed. This study employs the MLLM for the GUI testing, which is more effective through aligning the GUI screenshot with text information for the vision prompt.
@@ VII Conclusion
With the advancement of LLM, there have been tremendous applications of LLMs in software engineering-related tasks such as test case generation with great performance improvement. Subsequently, MLLMs emerge and are applied in image-text reasoning, video generation, etc., further expanding our understanding of the capabilities of LLMs. this paper proposes Trident, a novel vision-driven, multi-agent collaborative automated GUI testing approach for detecting non-crash functional bugs. It comprises three agents: Explorer, Monitor, and Detector. This work contributes as a stepping stone for future exploration in software engineering, offering insights and guidance on the potential applications of MLLM.
In the future, we will further improve this approach for better performance. Meanwhile, beyond Android apps, we will customize and apply our approach to other platforms such as web, desktop and automotive OS for software testing.